OpenStudy (anonymous):

Which of the following constants can be added to x2 - 3x to form a perfect square trinomial?

OpenStudy (anonymous):

9/4=2,25

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

How did you figure it out?

OpenStudy (mathstudent55):

Find the coefficient of the x term Divide it by 2. Then square it.

OpenStudy (mathstudent55):

Find the coefficient of the x term: -3 Divide it by 2: \(-\dfrac{3}{2} \) Then square it: \(\dfrac{9}{4} \)

OpenStudy (accessdenied):

A perfect square trinomial factors into this form: \( (x - h)^2 \) If we expand this form out, we get the perfect square trinomial. \( x^2 - 2h x + h^2 \) We may compare the coefficients between these two formats: \(x^2 = x^2 \) \( \color{blue}{-2h} x = \color{blue}{-3}x \) ==> 2h = 3 ==> h = 3/2 --> h^2 = (3/2)^2 goes below. \( \color{blue}{h^2} = \color{blue}{\text{constant added}} \) h^2 = constant This is where the mechanical process comes from: We always would set -2h ( from the expanded form of x^2 - 2hx + h^2) equal to the linear coefficient, and h^2 is always the constant addition. The process is taken to be (divide linear coefficient by 2) (square the result to find constant for perfect square)

OpenStudy (anonymous):

Thank you so much!! You helped a lot.

OpenStudy (accessdenied):

you're welcome!